### v1

# Interaction of water surface waves with periodic and quasiperiodic cylinder arrays

Abstract
Inspired by transformation optics and photonic crystals, this paper presents a computational investigation into the interaction between water surface waves and array waveguides of cylinders with multiple previously unexplored lattice geometries, including, for the first time, quasiperiodic geometries. Extending beyond conventional square and hexagonal periodic arrays, transformation optics has opened up entirely new opportunities to investigate water wave propagation through arrays based on quasiperiodic lattices, and quasiperiodically arranged vacancy defects. Using the linear potential flow open-source code Capytaine, missing element and -scaled Fibonacci square lattices, the Penrose lattice, hexagonal lattice and Amman-Beenker lattice are investigated. The existence of band gaps for all arrays is observed. An hexagonal lattice with vacancy defects transmits the least energy. Bragg diffraction consistent with rotational symmetry is observed from all arrays. Waves will distort significantly to achieve resonance with arrays, supporting transformation-based waveguides. The possible uses include adaptation to more versatile waveguides with applications such as offshore renewable energy and coastal defence.
I Introduction
Control of water wave propagation is an essential capability for multiple applications in coastal and offshore engineering. The power in surface waves can pound cliffs and beaches – contributing to the increasing global problem of coastal erosion – or it can be harnessed to provide energy to help break society’s reliance on fossil fuels. Improving our ability to manage and harness this energy is a problem of paramount importance, and has been the focus of much work over recent years [41, 39, 40, 36, 34, 15, 16, 17, 33, 38, 23]. Indeed, free-surface water waves (disturbances of water) have been comprehensively studied analytically [24], numerically [19], and through observations both in controlled physical experiments and field measurements [6, 7]. It has been shown that water surface waves share many of the same features as waves in other areas of physics (including fundamental processes such as reflection, refraction, diffraction, etc.). However, water waves also exhibit unique behaviour including shoaling (change of wave height) and breaking (collapse of waves) when they propagate into progressively shallower water. While the increasing emphasis on extracting energy from ocean waves and coastal protection has expanded our knowledge and understanding, new research is still needed, particularly to improve our understanding of complex wave interactions and the exploitation of wave energy devices.
One avenue of exploration over the last few decades has been the investigation of the interactions between water surface waves and arrays of cylinders, particularly in the context of offshore engineering. This has received renewed interest with the recent focus on offshore renewable energy and alternative coastal protection schemes. Interestingly, but not surprisingly, it has been shown that the problem of large arrays is analogous to electron propagation through crystals, leading to the application of the Bloch theorem to infinite and finite arrays of cylinders, and consequently the derivation of the band structure of these arrays [25]. Similarly, Ha et al. generated transmission spectra of finite square and hexagonal arrays for wave propagation along high-symmetry lattice directions, and found complete band gaps in each case [14].
In addition to periodic lattices built on crystalline symmetries, quasiperiodic lattices may also be constructed that have no translational symmetry but higher-order rotational symmetries, such as the pentagonal Penrose tiling [28] and the octagonal Ammann-Beenker tiling [3]. Quasicrystals (quasiperiodic crystals) based on these lattices are the appropriate analogue of periodic crystals [31, 21]. The higher orders of rotational symmetry available in quasiperiodic structures are attractive because they offer the same behaviours that periodic structures do, such as band gaps, but with greater isotropy and hence less angular selectivity [30].
The similarities observed in electron and water surface wave propagation have also motivated analogies to other wave physics, such as transformation optics. Transformation optics is a relatively new field devoted to tailoring the propagation of electromagnetic waves which has had remarkable breakthroughs for superlenses and invisibility devices [27]. Particular success has been found in constructing photonic crystals: metamaterials with structures on length scales similar to the wavelength of the radiation. Photonic crystals and photonic quasicrystals have both been shown to exert deep influence on electromagnetic wave propagation [18, 13]. In this case, modifying the permittivity and the permeability of the propagation medium manipulates the propagation of the waves. A similar field is transformation acoustics, in which the Young’s modulus of the propagation medium is modified, with similarly useful results [8]. The analogous properties which affect propagation of water waves are gravity and depth. Quasicrystalline metamaterials have been shown to have fractal transmission spectra with several large gaps [11]. Despite the clear analogies between transformation optics and water wave propagation, such lattice arrangements of cylinder arrays and scatterers are yet to be fully investigated for water waves.
With respect to water surface waves, and in the long-wavelength regime (, where is the wavelength and is a characteristic crystal dimension), it has been shown that the presence of an array of vertical cylinders alters the refractive index via modification of the effective gravity within the array, allowing construction of lenses and prisms [15]. This approach has also been shown experimentally to result in superlensing [17]. Moreover, analogous to the manipulation of permittivity/permeability in photonic materials, manipulating the depth via modification of either the floor or the free surface can lead to the full range of transformation optics effects [37, 29]. For example, the refractive index can be varied continuously by applying a gradient to the floor [35], or a band gap can develop via interaction with a periodically drilled floor [16]. However, there has been limited investigation into the exploitation of rotational symmetries within the lattice array structures.
Here, inspired by the similarities with transformation optics, we explore the interaction between monochromatic water surface waves and small () arrays of vertical surface-piercing cylinders. These cylinders are arranged in translational and rotational symmetries previously uninvestigated and, in particular, we measure their capacity to allow or stop wave propagation in the near-Bragg-resonance regime (). Our focus is on the detailed structure of the arrays, which include periodic, quasiperiodic, and defected periodic geometries with the resulting effects on water surface wave propagation.
The paper is structured as follows. First, we introduce the arrays that are investigated, and characterise them in terms of their symmetries. Then, we describe the wave modelling and numerical setup using the open-source software Capytaine [2, 4], before presenting the results from the simulations. We finish with a discussion on future work and conclusions.
II Lattices
We differentiate between arrays and lattices as follows: an array is a realization of a geometric lattice consisting of an arrangement of cylindrical free-surface-piercing scatterers.
Our interest lies in exploring the wave propagation properties of arrays with varying rotational symmetries and periodic/quasiperiodic structures. The designs investigated are shown in Figure 1, encompassing archetypes and simple modifications thereof of 4–, 6–, 8–, and 10–fold rotational symmetries, where they are generally referred to by their -fold symmetry as f. The finer details of their structure and nomenclature are discussed below. In an effort to study the effects of geometry independently of other characteristics of the array, where possible, the most common nearest neighbour distance is fixed at m. Cylinders are placed at the lattice points (section III.2) to calculate the filling fraction (FF) defined as the area occupied by cylinders divided by the total area of the array (see Figure 2(b)). According to the work of Hu et al.[15], the refractive index is then . Number of elements, FF and are given for each array in Table 1. For the purposes of comparison both periodic and aperiodic arrays are investigated.
Array | elements | Filling fraction (FF) | Refractive index |
4f | 264 | 0.136 | 1.066 |
4f | 199 | 0.118 | 1.057 |
4fτ | 288 | 0.184 | 1.088 |
6f | 250 | 0.150 | 1.072 |
6f | 202 | 0.110 | 1.054 |
6f | 232 | 0.138 | 1.067 |
8f | 240 | 0.153 | 1.074 |
10f | 246 | 0.154 | 1.074 |
II.1 Periodic arrays
The two simplest classes of periodic lattice are 4–fold (square) and 6–fold (hexagonal), arrays based on which are shown in Figure 1 and labelled as 4f and 6f respectively, where the subscript s denotes ‘simple’. All continuous lattices in two-dimensional space must have 2–fold rotational symmetry.
II.2 Quasiperiodic arrays
The remaining arrays are quasiperiodic. Here, we discuss the structure and characteristics of their underlying lattices.
Missing-element square Fibonacci lattice
Array 4f is based on a square lattice with elements missing in an quasiperiodic fashion. The missing elements are chosen based on the binary Fibonacci sequence (LSLLSLS..), where the individual elements of the sequence are either 2a (S, short) or 3a (L, long).
-scaled square Fibonacci lattice
Array 4fτ is a square lattice of elements placed according to 2 perpendicular Fibonacci sequences, where is equal to the golden ratio . As there are times as many segments as segments in each (infinite) sequence, the in this case is , and the FF is accordingly high [22].
Hexagonal lattices based on the lattice
Array 6f is based on a periodic hexagonal lattice with elements removed to produce a quasiperiodic arrangement, similar to 4f. The lattice points decorate the vertices of the lattice identified by Coates et al. [10], where the short and long edges of the rhombic and hexagonal tiles are both set to 1 m [9]. Array 6f is also produced using the lattice with tile edges set to 1 m, but instead uses the centre points of the tiles.
Penrose lattice
Array 10f is formed from the vertices of a Penrose P3 lattice [28], which is perhaps the most well-known quasiperiodic geometry. The lattice is composed of two types of rhombi, with internal angles (thin) and (fat), arranged according to matching rules such as to give the tiling an overall 10–fold symmetry, although it is often described as pentagonal.. The rhombus edge length is chosen to be 1 m. This again leads to elements being separated by less than 1 m, here, along the short diagonal of the thin rhombi.
III Methods
III.1 Numerical domain
The array is placed at the centre of a 40 m 40 m square grid, with a grid resolution of 512 512 and a resultant cell size of 7.8 7.8 cm. The waves are incident from the left hand side. They impinge on the array, are scattered, and we observe the steady state response of the water wave properties in the domain.
The water is set to infinite depth so that the waves can be considered deep water waves. The arrays themselves are composed of fixed 0.2 m radius, 10 m long, 13-gon prism approximations to cylinders, with 5 m submerged. As noted, the most common separation present in the array is set to 1 m. Wave heights are scaled from unity in Capytaine: any wave height or motion amplitude can be retrieved by multiplying the result by the desired value [2], and the results are valid in the linear regime.
III.2 Water wave modelling with Capytaine
The interaction between the waves and the arrays is simulated herein using the Capytaine open-source software (Ancellin and Dias [2] and Babarit and Delhommeau [4]). Capytaine is a Python-based boundary element method (BEM) solver for linear potential flow in water waves [1]. Based on previous code known as NEMOH [4], the linear potential flow approach has been widely used including multi-frequency and multi-direction wave loads on wind turbine platforms [20]. Starting from the assumptions of inviscid, irrotational and incompressible flow, the linear potential flow theory solves the problem in the frequency domain and is able to predict the radiation and diffraction processes. As a potential flow solver, the approach can be used to predict waves until they become near nonlinear and eventually break, at which point alternative (and computationally more expensive) approaches are required, such as full Navier-Stokes solvers ([26] and [12]). The predictions herein are only linear, so as the frequency increases, the maximum wave height for which they are applicable decreases. According to the limit for strict linearity , our waves are perfectly modelled only for a height between 4.3 cm at 3.0 rad s -1 and 6 mm at 8.0 rad s -1.
The calculations were performed on the Jeremiah Horrocks Institute High Throughput Cluster, which comprises 13 Dell R650XS nodes each with 48 physical cores and 256 GB RAM running Oracle Linux 8.9 and Slurm 20.11.9.
The arrays are constructed using Capytaine’s internal routines. One cylinder is placed, then duplicated to create the array. Tridecagonal prisms are used as they maximise usage of available RAM.
Capytaine can provide many kinds of output. Here we use the water free surface, provided as separate grids of the real and imaginary components of the solution to the wave equation. We use MATLAB to present and perform calculations on these datasets. An example dataset is shown in Figure 3, showing the interaction between waves and the 6f array. The two components of the free surface are shown in panels (a,c). All data in this manuscript presented with this colourmap are normalized to the range of the individual dataset.
We generate a reference wavefield of appropriate frequency and phase to represent the incoming waves and subtract it from the real component to give panel (b). This treatment aids in identification of scattered waves, in particular their direction. However, if there is any significant interaction between the waves and the array (i.e., if a phase difference is introduced, and/or the amplitude is altered), the area to the right of the array is dominated by the subtracted wavefield.
In panel (d) we add the real and imaginary components squared, which gives the intensity of the wave energy. Data presented with this colourmap always represent intensity, and are normalized to the range 0-5, with 1 being the same intensity as the incoming waves.
In panel (e), a polar plot shows the angular distribution function (ADF). This function is generated by taking a profile from the centre to the periphery of the intensity map, at the denoted angle, and plotting the mean value of this profile as a point on the red curve. For profiles of either the real or imaginary components, it would be appropriate to use the root mean square; as we use profiles of the intensity, there is no inherent periodicity and so the mean is appropriate. The influence of the varying length of the profiles in the non-circular data is minimized by use of the mean rather than the integral. In the example given, this shows strong intensity in the directions of scattered waves, and zero intensity in the direction of propagation after interaction with the array. The ADF calculation includes the area inside the array.
This system has mirror symmetry about the horizontal axis. The apparent asymmetry in the results, particularly noticeable in the ADF, is due to aliasing between the array and the simulation cell grid, and would be reduced for a higher-density grid. Another non-mirror-symmetric element, shown in Figure 2, is the 13-gon prism used to represent a cylinder, though the effect of this is negligible.
IV Results and discussion
Here, we present and discuss the results of our simulations in terms of relevant wave phenomena: diffraction, blocking, and resonance, before directly comparing the periodic and quasiperiodic arrays in terms of their performance.
IV.1 Diffraction
In general, all of the arrays produce strong diffraction when the Bragg condition is satisfied by one or more separations in the array. The diffracted beams are directional and diffuse, which is consistent with the small number of scatterers in the arrays.
We use the Bragg law to extract the apparent interplanar distance of the arrays when the wavefront is normal to the array. Close to the Bragg angles for a given lattice, constructive interference may occur for a range of frequencies, with the diffracted beam sweeping a concomitant range of angles. Therefore, it is appropriate to use the Bragg law to extract only when a lattice has one or more easily identified directions of symmetry. Figure 4 shows this analysis for the 4f, 6f, 8f, and 10f arrays, where a black arrow indicates the incoming and diffracted wave direction and angle associated with the rotational symmetry of the array (). For the simple square array, is found to be close to the expected value of , and for the simple hexagonal array, it is close to twice the expected value of , as our frequency range is too low to observe primary Bragg reflection. The deviation from the expected values is consistent with the finite size of the array and the finite radius of the scatterers. For example, for array 4f, scattering is evident from the entire range of lattices between those defined by the innermost points and outermost points of the cylinders in the array. In this array, the minimum number of planes at 45∘ is 11. This gives an approximate % error from the limits of , given by (.
The 4f, 6f, and 6f arrays show the same broad behaviour as the simple structures with this analysis. The 8f and 10f arrays have well-defined scattering angles, which permits the extraction of an apparent ‘interplanar’ distance, shown in the figure as .
The -scaled 4–fold array 4fτ, not shown here, has ambiguous behaviour: although it has 4–fold symmetry, it never strongly diffracts from a set of planes at 45∘ as expected. However, weak diffraction at multiple wavelengths yields values for of 1.00 m and 1.39 m.
IV.2 Blocking
The capacity of an array to remove intensity from an area behind the array versus the angular frequency of the incoming waves , was measured to give ‘blocking’ curves, which are presented in Figure 5 for each array. Blocking dips correlate with complete or incomplete band gaps. Using the ADF described above to compile this data presents two problems: firstly, it is strongly directionally selective; secondly, its calculation includes data inside the array. These problems limit its applicability to the question of how effective a given array is at blocking waves, so we take another approach. In the inset in Figure 5, we indicate a region of interest (ROI) to the right of the array. All datasets in the figure are generated using this ROI, with the mean value of intensity (as defined above) within the ROI used to represent the capacity of an array to remove intensity from the ROI.
A numerical summary of these results is compiled in Table 2, where we show the maximum blocking % for any frequency, and the total blocking observed over the full frequency range. Earlier work by Ha et al. using the multiple-scattering method – in which diffracted waves are linked to the incoming wave and represented by Fourier–Bessel expansions to derive the band structure of similar arrays – shows the existence of complete band gaps at for a square array and for an hexagonal array [14]. Our results show blocking dips starting at lower frequencies and bounded by these values at the upper end.
These results indicate that a specified periodic array can constitute an effective strategy for blocking water wave propagation for a given range of angular frequencies.
Arrays 4f and 6f, which have quasiperiodicity defined by vacancy defects, as opposed to intrinsic geometry, show effective broadband blocking, duplicating the single broad dip exhibited by 4f and 6f in addition to increased blocking at higher frequencies.
For the quasiperiodic arrays based on or higher rotational symmetries, the blocking curves are characterized by more dips, that individually are narrower than those for the periodic arrays. The behaviour of 4fτ is strikingly similar to 8f, in terms of the locations of the blocking dips. The action of the arrays over a wide angular range indicates that quasiperiodic lattices could contribute to broadband blocking strategies in addition to their rotational isotropy (which we do not investigate here).
The overall relative behaviour of the arrays according to their geometry is analogous to the situation of diffraction from crystals. Quasicrystals, due to their quasiperiodicity, have a diffraction pattern of infinite density. In practice, most of the diffraction peaks are too dim to observe, so a discrete diffraction pattern is observed. This pattern is of lower intensity than that from a periodic crystal, in which the scattering planes contributing to a particular low-index peak are far more numerous than those for a quasicrystal. Our quasiperiodic arrays 4fτ, 6f, 8f and 10f show analogous behaviour in their multiple blocking dips which each are smaller than those for the periodic arrays. This behaviour is reminiscent of the fractal transmission structure exhibited by quasicrystalline metamaterials [42, 11].
The difference between the 6f and 6f blocking curves is notable, as they are based on the same underlying lattice but with a different basis location. There therefore seem to be at least two independent components to the blocking behaviour: one dependent on standing waves in a periodic lattice, after the Bloch theorem, and one dependent on the underlying quasiperiodic ordering. In the blocking from 6f the curve is dominated by the component from the periodic array, whereas in the blocking from 6f, this component is entirely absent, revealing a blocking curve with the same characteristics as the other quasiperiodic arrays.
Maximum blocking | Maximum | Total blocking over | |
Array | per frequency (%) | (rad s | frequency range (%) |
4f | 99.2 | 5.6 [1.02] | 31.6 |
4f | 97.6 | 5.6 [1.02] | 32.9 |
4fτ | 84.2 | 6.0 | 23.4 |
6f | 98.8 | 6.8 [1.50] | 33.4 |
6f | 96.4 | 6.6 [1.41] | 36.5 |
6f | 54.8 | 6.3 | 15.6 |
8f | 70.1 | 6.1 | 15.4 |
10f | 68.0 | 6.3 | 16.7 |
IV.3 Resonance
Transmission through the array depends on resonance between some characteristic length in the array and the surface wavelength. Various resonances were observed across the arrays, corresponding with effective transmission. Some examples are shown in Figure 6, in which the intensity maps for the 4f array are reproduced for certain values of . Resonances are indicated, as are datasets far from resonance. These follow a notation in which the numerator is the number of lattice parameters and the denominator is the number of wavelengths. As also shown, when far from resonance, reflection of the waves dominates, forming strong standing wave patterns on the leading side of the arrays and a concomitant shadow on the trailing side.
The angular frequency is an input variable, and runs from 3.0 – 8.0 rad s-1 in increments of 0.1. The incident wavelength is calculated from . The wavelength inside the array is given by , where is the refractive index calculated from the filling fraction of the array [14]. The actual wavelength is expressed via the resonance ratio, given that the lattice parameter of the array is 1 m.
The difference between actual resonant wavelength and the expected wavelength inside the array is used to calculate the fractional change of wavelength , per wave, that is needed to achieve resonance with the array:
For array 4f, resonance is strongly correlated with transmission, and, to achieve resonance, the wavelength inside the array may be distorted by up to around . Higher order resonances are more distorted in our data, but that may be due to the angular frequency resolution, as the 0.1 rad s-1 steps become more significant at higher angular frequencies. The large degree of distortion achieved to gain resonance is strongly suggestive that the waves will follow such a resonance even if the array undergoes a moderate transformation. We leave this possibility for future work.
The corresponding intensity maps for array 4f, based on a square lattice with elements removed in a Fibonacci sequence, are shown in the lower half of Figure 6 for comparison. The resonances observed in 4f are strongly disrupted in 4f, resulting in reduced transmission. This effect becomes more pronounced with increasing , resulting in the increased blocking of array 4f over the entire frequency range.
V Link to transformation optics
As mentioned in the Introduction, this work has been partly inspired by the rise in transformation-based approaches to wave physics. Generally, this involves transformation (modification of) the constants in a wave propagation equation; permittivity and permeability for optics, Young’s modulus for acoustics, and gravity and depth for water surface waves. Modifying the density of a lattice of elements of given permeability/permittivity [Young’s modulus] causes optical [acoustic] waves to propagate in the direction of the density modifications. The situation for water surface waves is essentially the same. Some wave equations are summarized in Table 3.
A water surface waveguide array can modify the effective gravity and effective depth in the region of the array. This is in contrast to modifications of bathymetry, which allow control only over the effective depth, and which, in the context of coastal defence, are expensive and temporary.
Here we have presented the results of linear wave propagation through several array types. We have shown the existence of band gaps for all the arrays, strongly indicating that this is a viable strategy for blocking and/or reflecting wave energy. However, the relevance of our results in the context of transformation optics comes rather from the results shown in Figure 6; specifically, the resonance-led propagation modes. These provide insight into the potential engineering of resonant array waveguides to guide propagation of, rather than block, water surface waves. From simple inspection of Figure 6, it is clear that the waves are in some kind of resonance for a greater range of frequencies than is blocked by band gaps.
optics | acoustics | water surface waves |
VI Conclusions
We have used the open-source Capytaine software to investigate the behaviour of water surface wave array waveguides for a range of array geometries, with the number of array elements ranging from 199-288. Using this approach, we have successfully replicated an earlier result that finds band gaps for some of these array types [14]. Our methodology allows us to directly observe many wave phenomena in real space, for example Bragg diffraction, refraction and resonance. Significantly, our use of simple archetypes of periodic and quasiperiodic lattices demonstrates the potential of this approach to investigate any kind of array waveguide within the linear water wave regime.
A particularly striking result is the degree of reflection that is achieved by many of these geometries. Nearly 100% reflection is observed in several cases, strongly supporting the adoption of such arrangements in coastal defence strategies.
Interestingly, transmission through the waveguides is heavily influenced by resonances between the waves and the array geometry. These resonances can be easily disrupted by, e.g., removing array elements, further enhancing the ability of array waveguides to control wave propagation. The resonances suggest that a transformation-optics-inspired approach of modifying the arrays could be harnessed to aid propagation in preferred directions.
Of the arrays tested, the periodic arrays and two quasiperiodic arrays are characterised by a single lattice parameter. The blocking curves of these arrays are characterised by two large blocking dips.
The quasiperiodic arrays based on the golden ratio or on higher-order rotational symmetries generated blocking curves characterised by several blocking dips that are smaller than those for periodic arrays.
The array with the most effective blocking over the frequency range is hexagonal with quasiperiodically located defects. This array also has the lowest filling fraction of those studied.
In this work, we have normalised the arrays by using the same array element dimensions. To extract the detailed behaviour of geometry in isolation from such confounding factors as array density, it may be instructive to perform the study normalised to other factors, such as the array refractive index, which can be modified by changing element radius and thus filling fraction.
Finally, the use of full non-linear potential flow calculations in future applications of this approach would allow investigations of energy transfer between different modes.
VII Acknowledgments
JAS would like to acknowledge useful discussions with P. Moriarty, R. Lifshitz, M. Ancellin and B. A. Patterson. The figures in this manuscript were produced using Inkscape, Plot2 and Veusz. The data were analysed using MATLAB and Fiji.
Funding: SC was supported by EPSRC grant EP/X011984/1. The computational work was supported by the Jeremiah Horrocks Institute.